# HeightFormer: Explicit Height Modeling without Extra Data for Camera-only 3D Object Detection in Bird’s Eye View

### v2

@@ I Introduction
Vision-based Bird’s Eye View (BEV) representation[1, 2, 3, 4, 5] is an emerging perception formulation for autonomous driving. It transforms and maps the information from the image space to a unified 3D BEV space, which can be used for various perception tasks like 3D object detection and BEV map segmentation. Moreover, the unified BEV space can directly fuse other modalities like LiDAR without any cost, which is of great scalability.
@@
Motivated by the above observations, we propose to explicitly add and model extra conditions to realize better 2D to 3D mapping. Similarly, some works[8, 9] propose to directly learn depth as the extra condition with depth pre-training or LiDAR information. Different from using depth, we explicitly model the height condition in the mapping for the following reasons. First, we prove that height in the BEV space is equivalent to depth in the image space for the 2D to 3D mapping problem. Both ways can provide equivalent conditions to resolve the problem of mapping. In this way, we can realize well-defined one-to-one mapping between 2D and 3D. Second, the height information in the BEV space can be retrieved from the BEV annotations without any other data modalities like LiDAR, while depth condition needs extra pre-training or LiDAR. In this work, we use the height information from the object’s 3D bounding box, which can be directly accessed from the ground truth. Third, the modeling in height can fit arbitrary camera rigs and types. For example, on NuScenes[10], the focal length of the backward camera is different from other cameras, resulting in different depth estimation patterns. In other words, different depth estimation network is needed for different cameras. While for the height condition, no matter which kind of camera configuration is used, it is processed with the same pattern in the BEV space. In this way, the height condition is more robust and flexible.
In this work, we propose a network that explicitly models height in the BEV space, which fulfills the condition needed for 2D to 3D mapping, termed as HeightFormer. Moreover, based on the height modeling, self-recursive height predictors are proposed to introduce the uncertainty of heights and segmentation maps which are used in the BEV query mask mechanism to produce high-precision detection results. In summary, the main contributions of our work are summarized as follows: 1) We give theoretical proof of the equivalence between height-based methods in BEV and depth-based methods in images, which is the basis of our work. The proof also demonstrates the feasibility of detection in the BEV space generated with predicted heights. 2) We propose to explicitly model heights in the BEV space without extra LiDAR supervision. A self-recursive predictor is proposed to model heights and a corresponding segmentation-based query mask is designed to handle positions whose heights cannot be defined. 3) Experiments on NuScenes[10] show that the proposed HeightFormer achieves the SOTA performance. Extensive quantitative and qualitative results show that it is feasible and effective to model heights in the BEV space and construct the BEV representation with predicted heights. The generalization analysis also shows that the proposed method can be applied to different methods, as a plugin and as compensation for depth modeling.
@@ II Related work
Depth estimation is essential for 3D object detection. Due to the high similarity between depth estimation and height estimation, we learn from the experiences in depth estimation. Early works focus on geometry-based methods for stereo images[29, 30, 31, 32]. In the monocular situation, there are two mainstream ways to estimate depth, which are direct regression by a well-designed network[33, 34, 35, 36, 37, 38] and geometry depth derived from the pinhole imaging model[39].
@@
Compared to depth modeling, constructing BEV features via height modeling is adopted by fewer works. BEVFormer[25] and PolarFormer[46] obtain BEV queries by sampling 3D points around fixed heights and assigning to them different attention weights, which is an implicit way of height modeling. BEVHeight[47] proposes to model object heights by predicting heights for image pixels and then lifts image features to BEV space via geometric transformation, which relies on the high camera installation and works better in roadside situations.
Different from previous works, we explicitly estimate heights in BEV space and focus on car-side situations, in which cameras are mounted on the car roof and occlusion is common. This work will also prove the equivalence between the height-based solution and the depth-based solution for constructing the BEV representation.
@@ III-A Theoretical equivalence of depth and height
In this part, we give proof of the equivalence between depth modeling and height modeling in 2D to 3D mapping, as shown in fig. 2. Meanwhile, the proof shows the feasibility of constructing high-quality BEV features by modeling heights accurately.
@@ Conclusion
According to eq. 4, eq. 7, and assumption eq. 2, for the given feature misplacement error $\epsilon$, a depth prediction error of $\delta_{d,max}$ is equivalent to a height prediction error of $\delta_{y,max}$, because they are both proportional to $\epsilon$. As a result, considering two models, one models the height and the other models the depth. Their height or depth estimation error is proportional, and they yield features with similar misplacement error. Therefore, in terms of error modeling, they are considered equivalent.
@@ III-D Network design
Temporal fusion modules[25] fuse information from previous frames with Self Attention. In this module, the query is current BEV query, and the value is the concatenated (previous and current) BEV query. DETR-style decoders are adopted to detect objects based on BEV features.
@@ IV-B Ablation study
We first replace the height predictors with ground truth heights to evaluate whether accurate height estimation could improve the performance of detection, and we call this method HeightFormer-gt. For BEV girds containing objects, predicted heights are replaced with ground truth heights. For others, predicted heights are replaced with $y_{xz}=0.0$ and $h{xz}=1.0$, which degrades into the situation of BEVFormer. Results are shown in table I.
@@
$*{}^{*}$ 0.327 0.282
$*{}^{*}$ depth 0.515 0.470
@@
As shown in table I, HeightFormer-gt exceeds BEVFormer by 0.208 in NDS and 0.374 in mAP, which can be considered as the performance upper bound. If heights could be accurately modeled, there would be a giant leap in performance. In the meanwhile, Lift-Splat[6] in table I is a method that constructs the BEV space with depth modeling and according to BEVDepth[9] introducing ground truth depth improves mAP of the Lift-Splat style detector by 0.188. The experiments in table I show that there is much room for improvement whether in depth modeling or height modeling.
Besides, it should be noticed that the two types of methods, LSS and BEVFormer, cannot be compared directly as they have different network architectures. We list these results here for showing the potential of height-based methods.
@@
We present the vanilla HeightFormer which turns fixed reference points into adaptive reference points. The adaptive reference points are generated with predicted anchor heights. In this way, height information is explicitly learned by the height predictor. To show the effectiveness of explicit height modeling, it is compared with BEVFormer which encodes height information into attention weights of spatial cross-attention in an implicit way. The vanilla HeightFormer simply has a multilayer perceptron as the height predictor. Results are shown in table II. Explicit height modeling brings a gain of 0.5 percentage points in both NDS and mAP.
TABLE II: Ablation study on height modeling. *: Vanilla HeightFormer which predicts heights in a standalone way.
@@ IV-C LiDAR supervision
With the extra supervision of LiDAR information, the performance is improved by 0.3 percentage points. The ablation study is shown in table V.
@@
We can note the extra LiDAR supervision brings little improvement to the performance of the proposed method compared with the improvement brought by explicit height modeling. A potential reason is that LiDAR points are sparse and occluded positions are not covered. This might yield inconsistent supervision information. This is different from the situation of depth-modeling, in which most pixels can be paired with one or more LiDAR points. However, the advantage of our method is that, it is a cost-free approach for improving the performance and requires no extra data.
@@ V Conclusion and Limitation
$\left[\begin{array}[]{c}x_{gt}\\ y_{gt}\\ z_{gt}\end{array}\right]$ $=K^{-1}\left[\begin{array}[]{c}u_{gt}\cdot d_{gt}\\ v_{gt}\cdot d_{gt}\\ d_{gt}\end{array}\right],$ (25)
$K$ $=\left[\begin{array}[]{ccc}f_{x}&0&u_{0}\\ 0&f_{y}&v_{0}\\ 0&0&1\end{array}\right],$ (29)
where $(u_{gt},v_{gt})$ is the position of the object in the image. $d$ is the given depth at $(u_{gt},v_{gt})$. The feature at $(u_{gt},v_{gt})$ is gathered into the BEV grid at $(x,z)$. $K$ is the camera’s intrinsic matrix which does the transformation between the image frame and the BEV frame. [^4] Theoretically, there is a camera frame other than the BEV frame. Without losing generality, this manuscript does not distinguish between the camera frame and the BEV frame, because we assume that the extrinsic parameter matrix is an identity matrix. Expanding eq. 18, eq. 25 and eq. 29, and we get

### v3

@@ I Introduction
Vision-based Bird’s Eye View (BEV) representation[1, 2, 3, 4, 5, 6, 7] is an emerging perception formulation for autonomous driving. It transforms and maps the information from the image space to a unified 3D BEV space, which can be used for various perception tasks like 3D object detection and BEV map segmentation. Moreover, the unified BEV space can directly fuse other modalities like LiDAR without any cost, which is of great scalability.
@@
Motivated by the above observations, we propose to explicitly add and model extra conditions to realize better 2D to 3D mapping. Similarly, some works[10, 11] propose to directly learn depth as the extra condition with depth pre-training or LiDAR information. Different from using depth, we explicitly model the height condition in the mapping for the following reasons. First, we prove that height in the BEV space is equivalent to depth in the image space for the 2D to 3D mapping problem. Both ways can provide equivalent conditions to resolve the problem of mapping. In this way, we can realize well-defined one-to-one mapping between 2D and 3D. Second, the height information in the BEV space can be retrieved from the BEV annotations without any other data modalities like LiDAR, while depth condition needs extra pre-training or LiDAR. In this work, we use the height information from the object’s 3D bounding box, which can be directly accessed from the ground truth. Third, the modeling in height can fit arbitrary camera rigs and types. For example, on NuScenes[12], the focal length of the backward camera is different from other cameras, resulting in different depth estimation patterns. In other words, different depth estimation network is needed for different cameras. For the height condition, no matter which kind of camera configuration is used, it is processed with the same pattern in the BEV space. In this way, the height condition is more robust and flexible.
In this work, we propose a network that explicitly models height in the BEV space, which fulfills the condition needed for 2D to 3D mapping, termed as HeightFormer. Moreover, based on the height modeling, self-recursive height predictors are proposed to introduce the uncertainty of heights and segmentation maps which are used in the BEV query mask mechanism to produce high-precision detection results. In summary, the main contributions of our work are summarized as follows: 1) We give theoretical proof of the equivalence between height-based methods in BEV and depth-based methods in images, which is the basis of our work. The proof also demonstrates the feasibility of detection in the BEV space generated with predicted heights. 2) We propose to explicitly model heights in the BEV space without extra LiDAR supervision. A self-recursive predictor is proposed to model heights and a corresponding segmentation-based query mask is designed to handle positions whose heights cannot be defined. 3) Experiments on NuScenes[12] show that the proposed HeightFormer achieves the SOTA performance compared with models not using LiDAR. Extensive quantitative and qualitative results show that it is feasible and effective to model heights in the BEV space and construct the BEV representation with predicted heights. The generalization analysis also shows that the proposed method can be applied to different methods, as a plugin and as compensation for depth modeling.
@@ II Related work
Depth estimation is essential for 3D object detection. Due to the high similarity between depth estimation and height estimation, we learn from the experiences in depth estimation. Early works focus on geometry-based methods for stereo images[31, 32]. In the monocular situation, there are two mainstream ways to estimate depth, which are direct regression by a well-designed network[33, 34, 35, 36] and geometry depth derived from the pinhole imaging model[37].
@@
Compared to depth modeling, constructing BEV features via height modeling is adopted by fewer works. BEVFormer[27] and PolarFormer[44] obtain BEV queries by sampling 3D points around fixed heights and assigning to them different attention weights, which is an implicit way of height modeling. BEVHeight[45] proposes to model object heights by predicting heights for image pixels and then lifts image features to BEV space via geometric transformation, which relies on the high camera installation and does not work well in car-side situations.
Different from previous works, we explicitly estimate heights in BEV space and focus on car-side situations, in which cameras are mounted on the car roof and occlusion is common. Detailed comparison can be found in the Appendix. This work will also prove the equivalence between the height-based solution and the depth-based solution for constructing the BEV representation.
@@ III-A Theoretical equivalence of depth and height
In this part, we give proof of the equivalence between depth modeling and height modeling in 2D to 3D mapping, as shown in fig. 2. The proof is aimed at showing the feasibility of constructing high-quality BEV features by modeling heights accurately. Besides, at the end of this part, we analyze the advantages of height modeling.
@@ Conclusion
According to eq. 4, eq. 7, and assumption eq. 2, for the given feature misplacement error $\epsilon$, a depth prediction error of $\delta_{d,max}$ is equivalent to a height prediction error of $\delta_{y,max}$, because they are both proportional to $\epsilon$. As a result, considering two models, one models the height, and the other models the depth. Their height or depth estimation error is proportional, and they yield features with similar misplacement errors. Therefore, in terms of error modeling, they are considered equivalent.
@@ III-D Network design
Temporal fusion modules[27] fuse information from previous frames with Self Attention. In this module, the query is the current BEV query, and the value is the concatenated (previous and current) BEV query. DETR-style decoders are adopted to detect objects based on BEV features.
@@ IV-B Ablation study
We first replace the height predictors with ground truth heights to evaluate whether accurate height estimation could improve the performance of detection, and we call this method HeightFormer-gt. For BEV girds containing objects, predicted heights are replaced with ground truth heights. For others, predicted heights are replaced with $y_{xz}=0.5$ and $h_{xz}=1.0$ (normalized), which degrades into the situation of BEVFormer. Results are shown in table I.
@@
Lift-Splat - 0.327 0.282
Lift-Splat-gt depth 0.515 0.470
@@
As shown in table I, HeightFormer-gt exceeds BEVFormer by 0.208 in NDS and 0.374 in mAP, which can be considered as the performance upper bound. However, in reality, height or depth estimation error could be amplified when considering the final detection error. As a result, it is not feasible to achieve this theoretical upper bound. In the meanwhile, Lift-Splat[8] in table I is a method that constructs the BEV space with depth modeling and according to BEVDepth[11] introducing ground truth depth improves mAP of the Lift-Splat style detector by 0.188. The experiments in table I show that there is room for improvement whether in depth modeling or height modeling.
Besides, it should be noted that the two types of methods, LSS and BEVFormer, cannot be compared directly as they have different network architectures. We list these results here to show the potential of height-based methods.
@@
We present the vanilla HeightFormer which turns fixed reference points into adaptive reference points. The adaptive reference points are generated with predicted anchor heights. In this way, height information is explicitly learned by the height predictor. To show the effectiveness of explicit height modeling, it is compared with BEVFormer which encodes height information into attention weights of spatial cross-attention in an implicit way. The vanilla HeightFormer simply has a multilayer perceptron as the height predictor. Results are shown in table II. Explicit height modeling brings a gain of 0.5 percentage points in both NDS and mAP for the “base” model. For the “tiny” model, the improvements come to 1.8 and 1.1 percentage points.
[…]
@@ IV-C LiDAR supervision
With the extra supervision of LiDAR information, the performance is improved by 0.7 percentage points. The ablation study is shown in table V.
@@
We can note the extra LiDAR supervision brings little improvement to the performance of the proposed method compared with the improvement brought by explicit height modeling. A potential reason is that LiDAR points are sparse and occluded positions are not covered. This might yield inconsistent supervision information. This is different from the situation of depth modeling, in which most pixels can be paired with one or more LiDAR points. However, the advantage of our method is that it is a cost-free approach to improve performance and requires no extra data.
@@ V Conclusion and Limitation
$\left[\begin{array}[]{c}x_{gt}\\ y_{gt}\\ z_{gt}\end{array}\right]$ $=K^{-1}\left[\begin{array}[]{c}u_{gt}\cdot d_{gt}\\ v_{gt}\cdot d_{gt}\\ d_{gt}\end{array}\right]$ (25)
where $(u_{gt},v_{gt})$ is the position of the object in the image. $d$ is the given depth at $(u_{gt},v_{gt})$. The feature at $(u_{gt},v_{gt})$ is gathered into the BEV grid at $(x,z)$. $K$ is the camera’s intrinsic matrix which does the transformation between the image frame and the BEV frame as stated in eq. 1.444Theoretically, there is a camera frame other than the BEV frame. Without losing generality, this manuscript does not distinguish between the camera frame and the BEV frame, because we assume that the extrinsic parameter matrix is an identity matrix. Expanding eq. 18, eq. 25 and LABEL:eq:d4, and we get
@@
Comparison to BEVHeight
In this appendix, we provide a detailed comparison between the proposed method and BEVHeight[45], which also constructs BEV features via height modeling.
(a) BEVHeight
(b) HeightFormer
Figure 11: Differences between BEVHeight and HeightFormer
  1. 1.
    Modeling: BEVHeight models s for each image pixel, and then projects image features into the BEV space. This is LSS-style[8] modeling, and it can be turned into depth modeling by mapping height bins into depth bins. In contrast, HeightFormer adopts OFT-style[9] modeling. BEV features are gathered from image features by projecting reference points into images.
  2. 2.
    Scenario: BEVHeight is primarily designed for roadside 3D object detection tasks, whereas HeightFormer is tailored for vehicle-mounted scenarios. BEVHeight operates optimally with cameras with high installation, as highlighted in its paper. The proposed method has no such requirements as it models heights with BEV features which have historical information about objects.
  3. 3.
    Height: BEVHeight focuses on modeling surface heights relative to the flat ground, while HeightFormer characterizes both the center and the range of an object along the height axis.
  4. 4.
    Implementation: BEVHeight employs a strategy of dividing heights into discrete bins and performing classifications for each image pixel. In contrast, HeightFormer utilizes Laplacian priors to model heights in contiguous space and conducts height regression for each BEV grid.
[…]
Model NDS mAP
BEVFormer 0.403 0.288
BEVHeight 0.342 0.291
HeightFormer 0.421 0.299
Furthermore, we compare the performance. The table IX shows the comparative experiments with BEVHeight. Because BEVHeight does not use history frames and its code does not contain training configurations on NuScenes, we compare mAP only. The proposed method outperformed BEVHeight by 0.8 percentage points.